# An Empirical Study of Cross-Lingual Transfer Learning in Programming Languages

### v2

@@ Abstract
Large language models (LLMs) have become remarkably good at improving developer productivity for high-resource programming languages. These models use two kinds of data: large amounts of unlabeled code samples for pre-training and relatively smaller amounts of labeled code samples for fine-tuning or in-context learning. Unfortunately, many programming languages are low-resource, lacking labeled samples for most tasks and often even lacking unlabeled samples. Therefore, users of low-resource languages (e.g., legacy or new languages) miss out on the benefits of LLMs. Cross-lingual transfer uses data from a source language to improve model performance on a target language. It has been well-studied for natural languages, but has received little attention for programming languages. This paper reports extensive experiments on four tasks using a transformer-based LLM and 11 to 41 programming languages to explore the following questions. First, how well does cross-lingual transfer work for a given task across different language pairs. Second, given a task and target language, how should one choose a source language. Third, which characteristics of a language pair are predictive of transfer performance, and how does that depend on the given task. Our empirical study with 1,808 experiments reveals practical and scientific insights, such as Kotlin and JavaScript being the most transferable source languages and different tasks relying on substantially different features. Overall, we find that learning transfers well across several programming languages.
Figure 1: Overview. Consider an LLM pretrained on unlabeled code in multiple seen languages. Finetune on task-specific labeled samples from a source language. For RQ1, test performance on a target language. For RQ2, train a model that ranks transfer performance given features of a language pair. For RQ3, measure how important the features of language pairs are. Repeat for several languages and four tasks: Error Detection, Solution Domain Classification, Clone Detection, and Code Repair.
I Introduction
@@
Large language models (LLMs) leverage the naturalness of software [1] to achieve state-of-the-art performance on many software engineering tasks such as error detection, clone detection, and code translation [2, 3]. However, LLMs have thus far been trained and evaluated mainly on programming languages with vast amounts of openly available code rather than low-resource ones. Programming languages used for critical and strategic applications may have much code only in private repositories and are thus ignored. COBOL, which supports many critical financial applications and may have over 775 billion lines of code overall [4], is low-resource due to lack of open availability. Similarly with Fortran, which is used in high performance and scientific computing domains. Notwithstanding their popularity [5, 6], new languages such as Rust and Swift start out low-resource, by definition. Other languages such as R may be low-resource for licensing reasons [7]. Besides being underrepresented in training of LLMs, organizations often spend significant resources on training software developers to work with unfamiliar low-resource languages. This adds to the cost of maintaining such software and may also stifle innovation [8]. There is a need for state-of-the-art AI tools to augment the capabilities of developers for these low-resource languages. Indeed, LLMs are being used in tools to enhance developer productivity and can also help with migration or modernization of projects from one programming language to another by supporting code translation [9, 10].
@@
Recent works have shown the possibility of transfer learning: leveraging data from one programming language to compensate for the lack of data in the target programming language [11, 12, 13]. Another recent work has developed several similarity metrics to decide which high-resource programming language dataset can be used to augment fine-tuning data for some other language [14]. Prior work, however, has used at most 6 programming languages, none of which are truly low-resource. Further, the notion of similarity among programming languages is underexamined: more research is needed to make reliable claims. To overcome these shortcomings, we examine transferability from source to target language combinations for 4 tasks and for many more languages than prior work (11–41 depending on the task). The four tasks we focus on are clone detection, code repair, solution domain classification, and error detection. Fig. 1 gives a high-level overview of the methodology for our extensive empirical study. From left to right, we start with a pretrained LLM that has seen unlabeled code in several programming languages. Next, for each of four tasks, for each source language with a suitable amount of labeled training data for that task, we finetune the LLM. For each target language with a suitable amount of labeled test data for the task, we evaluate the finetuned LLM. For example, Fig. 1 illustrates this with a solution domain classification task, using C as the source language and Rust as the target language, which yields an F1 score of $0.48$. Doing this for four different tasks and all their source and target language combinations yields four different heat maps.
@@
Our first research question (RQ1) asks how learning transfers across programming languages in general, by exploring these heat maps. Our second research question (RQ2) asks whether one can predict the ranking of source languages in the heat maps from the previous experiment just using features of language pairs, by training a performance prediction model from the language pair features and the ground-truth target labels from the heat maps. For example, in Fig. 1, the performance prediction model for the solution domain classification task might rank C behind Kotlin as a source language for transfer to Rust. Given the cost of training LLMs, the performance prediction model is useful in directing data acquisition efforts and deciding how to spend compute resources. Finally, as our third research question (RQ3), we seek software-engineering insights into the characteristics of programming languages that affect transferability. To do this, we measure the importance of language pair features in the prediction model from the previous experiment.
@@
  • Evaluating the pairwise transferability for several programming languages (including low-resource ones) and four tasks using LLMs.
  • Developing a method to identify the best language to transfer from, for different target languages and tasks.
@@
We are planning to make our data generation, training, and analysis code available. One goal of this paper is to offer practical guidance to engineers who build LLM-driven software engineering tools towards more informed choices for data acquisition and modeling. A second goal of this paper is to help advance software engineering as a science by using the lens of transfer learning to shed light on programming language characteristics. Below are some results highlights:
  • Learning transfers well for all tasks. More specifically, cross-lingual learning transfers better than zero-shot.
  • Transfer learning depends on the target language. Java, Go, Dart, and TypeScript are among the best target languages. C++, Python, Vim, and COBOL are among the worst.
@@
  • We can predict which source language can perform well for a given task by considering a variety of features.
@@
  • Different tasks rely on substantially different features.
  • Overall the most important features are the overlap of keywords and names between the source and target languages.
@@
If we take the premise that software engineers will increasingly benefit from LLM support, then we hope this paper will help democratize such support into low-resource settings.
II Background and Related Work
Cross-lingual Transfer for Natural Languages
@@ II Background and Related Work
Lin et al. [15] use 4 tasks to study transfer among up to 60 natural languages and explore how important features of language pairs are for predicting how well learning transfers between them, finding that feature importances vary a lot across tasks. Lauscher et al. [16] use 5 tasks to study transfer among up to 15 languages and argue that while zero-shot performance can work for low-level tasks, higher-level tasks benefit from at least a few target-language shots. de Vries et al. [17] use 1 task (POS tagging) with 65 source and 105 target languages and study the effects of language families and writing systems, with Romanian emerging as a particularly good source language. Ahuja et al. [18] use 11 tasks with 1 source and varying numbers of target languages, confirming earlier findings that feature importances vary a lot across tasks. Our paper takes inspiration from these works, especially de Vries et al. [17], but differs by studying programming languages instead of natural languages, which have different tasks and different features of language pairs affecting transferability.
Cross-lingual Transfer for Programming Languages
@@
Zhou et al. [5] call for studying transfer, motivating with 1 task (code completion) and 2 languages (from Hack to Rust). Chen et al. [14] use 2 tasks (code summarization and search) and 6 languages to study transfer to Ruby, and propose picking the source language based on language similarity. Ahmed et al. [11] use 3 tasks to study transfer among 6 languages, demonstrating that due to the nature of their tasks, signals from identifiers are highly important for transferability. Yuan et al. [12] use 1 task (automated program repair) to study transfer among 5 languages, sequentially fine-tuning on multiple languages with innovative tricks to prevent catastrophic forgetting. Pian et al. [13] use 2 tasks (code summarization and completion) to study transfer among 4 languages, using meta-learning to improve a base learner. Our paper also focuses on programming languages, but considers more tasks and vastly more languages to explore conditions for effective transfer.
@@
Besides the above works, there is also work on learning for multiple programming languages. While such prior work has been instrumental in making studies like ours possible, it differs by not focusing on cross-lingual transfer. Our work benefits from datasets for multiple programming languages. CodeNet covers 55 languages [19]; XCodeEval covers 7 tasks with up to 17 languages [20]; and MultiPL-E supports 1 task (code generation) in 19 languages [21]. There are also models for multiple programming languages. TransCoder uses pre-training across 3 languages to help transpilers [10]. CodeT5 is pre-trained on 8 languages and 5 tasks [22]. StarCoder is pretrained on 86 languages [7], but e.g. does not include COBOL and evaluation is largely focused on Python. Unsupervised pretraining on multiple languages has become common, but transferability of supervised tasks has not yet been thoroughly studied; our paper addresses that gap.
III Experimental Setup
@@
As illustrated in Fig. 1, given a task, our experimental approach first finetunes a model for each source language individually and then tests each finetuned model on all target languages. We applied this approach to four tasks where the number of source languages varies from 6 to 22, leading to 58 finetuned models in total. For each source language, we finetune the model using one A100 GPU for 6-30 hrs depending on the task. Each model was then evaluated on 11 to 43 target languages producing 1,808 experiments. All results are presented in Figs. 5(a)–(d) and analyzed in Sec. IV.
This section presents methodological details.
@@
Figure 2: Transfer scores heat map. The figure shows scores for every combination of source and target language. Each column corresponds to a source language, with “zsh” showing zero-shot performance, i.e., without fine-tuning on any source language. Each row corresponds to a target language. The languages whose language-name label uses red font were seen during pre-training. Framed black boxes highlight the performance of a source language (column) on itself as the target language (row). The dendrograms show results of hierarchical clustering based on similarity of the performance vectors. The row and column order is also determined by the same clustering.
@@
Figure 3: Transfer scores heat map (cont.; see caption from 2(a))
@@
Figure 4: Transfer scores heat map (cont.; see caption from 2(a)).
@@
Figure 5: Transfer scores heat map (cont.; see caption from 2(a)).
Base Datasets
@@
The four tasks studied are derived from two multilingual code datasets: CodeNet [19] and XCodeEval [20]. CodeNet [19] consists of about 14M code samples for a total of 500M lines of code in 55 different programming languages, derived from submissions to online coding judge websites. The dataset comes with benchmarks for code classification and code similarity. XCodeEval [20] contains 25M samples in 11 programming languages from another different online coding judge website. This dataset comes with an execution-based evaluation framework and several different benchmarks: from classification (solution domain classification and error detection) to generation (program synthesis, automatic program repair, and code translation).
Tasks
@@
To study how learning transfers between programming languages, we explore two type of tasks: classification (error detection, solution domain classification, clone detection) and generation (code repair). For the classification tasks, we follow Lewis et al. [23] by predicting labels from the vocabulary of class labels based on the final decoder hidden state. We use BLEU score [^1] We chose BLEU score because it enables experiments with the larger number of languages in CodeNet [19], and because other prominent benchmarks for the code repair task also adopt this metric [24]. for evaluating the generative task and F1 score for evaluating the classification tasks.
@@
  • Error Detection: Given a code $C$ in language $L$, the task is binary classification of whether or not $C$ compiles (or can be loaded by an interpreter) without error.
@@
  • Solution Domain Classification: Given a code $C$ in language $L$, the task is multi-label classification of $C$ into a set of tags corresponding to potential algorithmic techniques required to write the solution (e.g., sorting, graphs).
@@
  • Clone Detection: Given two code samples $C_{1}$ and $C_{2}$ in language $L$, the task is binary classification of whether they are type-IV clones (semantically similar) [25]. To generate the dataset, we apply the following procedure to the CodeNet dataset. Given all combinations of solutions to all problems in language $L$, we identify positive samples (clones) as pairs of accepted solutions for the same problem and the others as negative examples. To balance the positive samples across problems, we ensure a ratio of 0.15 of positive samples across different languages.
@@
  • Code Repair: Given a buggy code $C$ in language $L$, the task is to generate the corresponding repaired code. We generate the dataset by using the following procedure on CodeNet solutions: Each sample is modified by sequentially inserting, removing, or replacing tokens of different types for fixed ratios for different token types.
Data Sampling
@@
Given variation in the size of datasets for different languages, we follow the sampling procedure of de Vries et al. [17] to select the number of training samples and identify potential source languages. We first finetune the model with datasets of different number of samples $N$={10K, 30K, 50K, 70K, 100K}. Depending on the number of training examples $N_{L}$ for a language $L$, we randomly upsample languages with $N<N_{L}$ or downsample languages with $N>N_{L}$. We filter the set of source languages for a given task based on the relative performance compared to a baseline model. We select sample sizes of 50K for solution domain classification, 70K for error detection, and 100K for clone detection and code repair. Fig. 8 in the supplementary material shows how the metric for each task vaires with the sample size.
Model
@@
Our experimental framework is based on CodeT5-base (220M parameters) leveraging the Hugging Face transformers library [26]. Code-T5 is an open-source model, pre-trained on eight programming languages (Ruby, JavaScript, Go, Python, Java, PHP, C, and C#). Due to its encoder-decoder nature, the model performs well on both code generation and code understanding tasks [22] like code repair, error detection, and clone detection, among others. This, along with its relatively small size, makes it a good fit for our empirical study, which requires 58 finetuned models with numerous inference runs each on four tasks.
IV Results and Discussion
This section addresses the following research questions.
RQ1:
How well does cross-lingual transfer work for a given task across different language pairs?
RQ2:
Given a task and target language, how should we pick the source language for best performance?
RQ3:
Which characteristics of a language pair are predictive of transfer performance, and how does that depend on the given task?
IV-A Transfer analysis

### v3

# Cross-lingual Transfer in Programming Languages: An Extensive Empirical Study

@@ Abstract
Large language models (LLMs) have achieved state-of-the-art performance in various software engineering tasks, including error detection, clone detection, and code translation, primarily leveraging high-resource programming languages like Python and Java. However, many critical languages, such as COBOL, as well as emerging languages, such as Rust and Swift, remain low-resource due to limited openly available code. This scarcity hampers the training and effectiveness of LLMs for these languages, increasing software maintenance costs and stifling innovation. Addressing this gap, we investigate the potential of transfer learning to enhance LLM performance on low-resource programming languages by leveraging data from high-resource counterparts. Our extensive empirical study evaluates transferability across 10 to 41 programming languages and five key tasks: code generation, clone detection, code repair, solution domain classification, and error detection. Additionally, we develop a performance prediction model to guess the best source languages for a given target and task, and analyze the features that influence transfer performance. We further replicate a representative subset of experiments with a larger model to test the generalizability of our conclusions to contemporary large-scale LLMs. Our findings demonstrate that cross-lingual transfer significantly outperforms zero-shot learning, with effectiveness varying based on both source and target languages. Languages such as Java and Go emerge as the best targets, while Kotlin and JavaScript are excellent sources. Furthermore, our model reliably predicts successful transfer sources by considering linguistic and dataset-specific features, offering practical guidance for data acquisition and model training. This work contributes to the development of LLM-driven tools for low-resource programming languages and provides insights into the characteristics that facilitate transfer across language pairs.
1 Introduction
Large language models (LLMs) leverage the fact that software code is often readable and repetitive (much like natural language (Hindle et al., 2016)) to achieve state-of-the-art performance on many software engineering tasks such as error detection, clone detection, and code translation (Lu et al., 2021). However, LLMs have thus far been trained and evaluated mainly on high-resource programming languages such as Python and Java with vast amounts of openly available code. Yet there are many low-resource programming languages that lack enough openly available code to train LLMs.
@@ 1 Introduction
COBOL, which supports many critical financial applications and may have over 775 billion lines of code overall (MicroFocus, 2022), is low-resource due to lack of open availability. Notwithstanding their popularity (Zhou et al., 2022; Coursera, 2024), new languages such as Rust and Swift start out low-resource, by definition. Other languages such as R are low-resource because most code in that language has a license that inhibits use for LLM training (Li et al., 2023). Besides being underrepresented in training of LLMs, organizations often spend significant resources on training software developers to work with unfamiliar low-resource languages. This adds to the cost of software maintenance and may also stifle innovation (Edwards & King, 2021). There is a need for state-of-the-art AI tools to enhance developer productivity for low-resource languages.
@@
Recent works have shown the possibility of transfer learning: leveraging data from one programming language to compensate for the lack of data in another (Ahmed & Devanbu, 2022; Yuan et al., 2022; Pian et al., 2023; Cassano et al., 2024). Another recent work has developed several similarity metrics to decide which high-resource programming language dataset can be used to augment fine-tuning data for a target language (Chen et al., 2022). Prior work, however, has used at most 6 programming languages, none of which are truly low-resource. Further, the notion of similarity among programming languages is underexamined: more research is needed to make reliable claims.
To overcome these shortcomings, we examine transferability from source to target language combinations for 5 tasks and for many more languages than prior work (10–41 depending on the task). We start by using training data for a task in a source language to fine-tune a small LLM (for 4 tasks) or to perform in-context learning using a larger LLM (for the 5th task). Then, we feed the respective LLM disjoint test data for the same task on a target language. Using this methodology, we evaluate the results across all source languages, target languages, and tasks. The source languages are, by definition, high-resource languages since we have enough data to train a model on them. Target languages include both high-resource and low-resource ones. The finetuning experiments focus on the four tasks clone detection, code repair, solution domain classification, and error detection. The in-context learning experiments use a fifth task, code generation.
Figure 1: Overview. Consider an LLM pretrained on unlabeled code in multiple seen languages. Finetune on task-specific labeled samples from a source language. For RQ1, test performance on a target language. For RQ2, train a model that ranks transfer performance given features of a language pair. For RQ3, measure how important the features of language pairs are. Repeat for several languages and four tasks.
Figure 1 summarizes the methodology for the finetuning part of our empirical study, which builds upon the work of de Vries et al. (2022) by exploring programming languages and multiple tasks. On the left, we start with a pretrained LLM that has seen unlabeled code in several programming languages. Next, for each of four tasks, for each source language with a suitable amount of labeled training data for that task, we finetune the LLM. For each target language with a suitable amount of labeled test data for the task, we evaluate the finetuned LLM. For example, Figure 1 illustrates this with a solution domain classification task, using C as the source language and Rust as the target language, which yields an F1 score of $0.48$. Doing this for four different tasks and all their source and target language combinations yields four different heat maps. The methodology using in-context learning on the fifth task is analogous and yields a fifth heat map.
This paper addresses three research questions (RQ1–RQ3).
RQ1: How well does cross-lingual transfer work for a given task across different language pairs? We answer this question directly by exploring the heat maps.
@@
RQ2: Given a task and target language, how should we pick the source language for best performance? We answer this question by training a performance prediction model from the language pair features and the ground-truth target labels from the heat maps, which predicts the ranking of source languages. For example, in Figure 1, the performance prediction model for the solution domain classification task might rank C behind Kotlin as a source language for transfer to Rust. Given the cost of training an LLM, the performance prediction model is useful in directing data acquisition efforts and deciding how to spend compute resources.
RQ3: Which characteristics of a language pair are predictive of transfer performance, and how does that depend on the given task? We answer this question by measuring the importance of language pair features in the prediction model from the previous experiment.
@@
  • Evaluating the pairwise transferability for several programming languages (including low-resource ones) and five tasks using language models.
  • Developing a method to predict the best language to transfer from, for different target languages and tasks.
@@
One goal of this paper is to offer practical guidance to engineers who build LLM-driven software engineering tools towards more informed choices for data acquisition and modeling. A second goal of this paper is to use the lens of transfer learning to shed light on programming language characteristics. Some highlights of our results are as follows:
  • Learning transfers well for all tasks and cross-lingual learning transfers better than zero-shot.
  • Transfer learning depends on the target language. Java and Go are the best target languages whereas C++ and Python are the worst.
@@
  • We can reliably predict which source languages will perform well on a given task by considering linguistic, synthetic, dataset-specific, and model-specific features.
@@
  • Different tasks rely on different features; for some, keywords and names are most important.
@@
Assuming software engineers increasingly benefit from LLM support, we believe this paper will help democratize such support into low-resource settings and help inspire more research in this area.
2 Related Work
Cross-lingual transfer for natural languages.
@@ 2 Related Work
Lin et al. (2019) use 4 tasks to study transfer among up to 60 natural languages and explore how important features of language pairs are for predicting how well learning transfers between them, finding that feature importances vary a lot across tasks. Lauscher et al. (2020) use 5 tasks to study transfer among up to 15 languages and argue that while zero-shot performance can work for low-level tasks, higher-level tasks benefit from at least a few target-language shots. de Vries et al. (2022) use 1 task (POS tagging) with 65 source and 105 target languages and study the effects of language families and writing systems, with Romanian emerging as a particularly good source language. Ahuja et al. (2022) use 11 tasks with 1 source and varying numbers of target languages, confirming earlier findings that feature importances vary a lot across tasks. Our paper takes inspiration from these works but differs by studying programming languages instead of natural languages, which have different tasks and different features affecting transferability.
Cross-lingual transfer for programming languages.
@@
Zhou et al. (2022) call for studying transfer, motivating with 1 task (code completion) and 2 languages (from Hack to Rust). Chen et al. (2022) use 2 tasks (code summarization and search) and 6 languages to study transfer to Ruby, and propose picking the source language based on language similarity. Ahmed & Devanbu (2022) use 3 tasks to study transfer among 6 languages, demonstrating that due to the nature of their tasks, signals from identifiers are highly important for transferability. Yuan et al. (2022) use 1 task (automated program repair) to study transfer among 5 languages, sequentially fine-tuning on multiple languages with innovative tricks to prevent catastrophic forgetting. Pian et al. (2023) use 2 tasks (code summarization and completion) to study transfer among 4 languages, using meta-learning to improve a base learner. Cassano et al. (2024) introduce MultiPL-T, a new semi-synthetic training dataset, and show that it improves transfer for 1 task (NL to code completion) across 19 languages. Our paper also focuses on programming languages, but considers more tasks and more languages to explore conditions for effective transfer.
@@
There is also work on learning across multiple programming languages that has been instrumental in making studies like ours possible, but it differs in that it does not focus on cross-lingual transfer. Unsupervised pretraining on multiple languages has become common, but transferability of supervised tasks has not yet been thoroughly studied; our paper addresses that gap. One could also imagine exploring transfer from multiple source languages to a low-resource target; we leave this to future work.
3 Experimental Setup
@@
Most of our experiments are based on finetuning a small model for four tasks as illustrated in Figure 1; in addition, there are also some experiments based on few-shot prompting a larger model without finetuning. For most experiments, given a task, our experimental approach first finetunes the CodeT5 (Wang et al., 2021) model for each source language individually and then tests each finetuned model on all target languages. We applied this approach to four tasks where the number of source languages varies from 6 to 22, leading to 58 finetuned models in total. For each source language, we finetune the model using one A100 GPU for 6 to 30 hours depending on the task. Each model was then evaluated on 11 to 43 target languages producing 1,808 experiments. All results are presented in Figures 5(a)–(d) and analyzed in Section 4. For the few-shot prompting experiments, we generated responses with the Llama 3.3 70B-Instruct model via TogetherAI, 2025 using a temperature of 0.8. Those results are presented in Figure 6.
@@
Figure 2: Transfer scores heat map. The figure shows scores for every combination of source and target language. Each column corresponds to a source language, with “Zero Shot” showing zero shot performance, i.e., without fine-tuning on any source language. Each row corresponds to a target language. The languages whose language-name label uses red font were seen during pre-training. Framed black boxes highlight the performance of a source language (column) on itself as the target language (row). The dendrograms show results of hierarchical clustering based on similarity of the performance vectors. The row and column order is also determined by the same clustering.
@@
Figure 3: (continued; see caption from 2(a))
@@
Figure 4: (continued; see caption from 2(a))
@@
Figure 5: (continued; see caption from 2(a))
(a) Code Generation: 10 source languages $\times$ 10 target languages. Metric: Pass@1 Score.
Figure 6: Few shot prompting transfer scores heat map. The figure shows scores for every combination of source and target language. Each column corresponds to a source language, with “Zero Shot” showing zero shot performance. Each row corresponds to a target language. Similarly, dendrograms show results of hierarchical clustering and framed black boxes highlight the performance of a source language on itself as the target language.
Base datasets.
@@
The five tasks studied are derived from three multilingual code datasets: CodeNet, XCodeEval, and MBXP. CodeNet (Puri et al., 2021) consists of about 14M code samples in 55 different programming languages, derived from submissions to online coding judge websites. The dataset comes with benchmarks for code classification and code similarity. XCodeEval (Khan et al., 2023) contains 25M samples in 11 programming languages from another different online coding judge website. This dataset comes with an execution-based evaluation framework and several different benchmarks: from classification (solution domain classification and error detection) to generation (program synthesis, automatic program repair, and code translation). MBXP (Athiwaratkun et al., 2023) is a multilingual extension of MBPP (Austin et al., 2021) to Java, JavaScript, TypeScript, Go, Ruby, Kotlin, PHP, C#, Scala, C++, Swift, and Perl. A companion code package is provided to perform execution in all supported languages. The dataset is designed for investigating the effectiveness of few-shot prompting, zero-shot translation, as well as the robustness to prompt perturbation. In our experiment, we utilize the few shot prompting setup.
Tasks.
@@
To study how learning transfers between programming languages, we explore two type of tasks: classification (error detection, solution domain classification, clone detection) and generation (code repair, code generation). For the classification tasks, we follow Lewis et al. (2020) by predicting class labels based on the final decoder hidden state. We use F1 score for evaluating the classification tasks, which is a widely accepted practice since it is good at dealing with class imbalance. We use BLEU score for evaluating the code repair task; while an execution-based metric would provide value, it was infeasible due to the large number of languages in CodeNet (Puri et al., 2021). (Other prominent benchmarks for the code repair task also adopt BLEU score (Lu et al., 2021).) Finally, we use execution-based Pass@1 score for evaluating the code generation task. The execution-based evaluation checks the functional correctness of the generated code by running it against the provided tests.
@@
Error Detection: Given a code $C$ in language $L$, do binary classification of whether $C$ compiles (or can be loaded by an interpreter) without error.
@@
Solution Domain Classification: Given a code $C$ in language $L$, do multi-label classification into a set of tags corresponding to algorithmic techniques required to write the solution (e.g., sorting, graphs).
@@
Clone Detection: Given two code samples $C_{1}$ and $C_{2}$ in language $L$, do binary classification of whether they are type-IV clones (semantically similar) (Roy & Cordy, 2007). We derived the dataset from CodeNet as follows. Given all combinations of solutions to all problems in language $L$, we identify positive samples (clones) as pairs of accepted solutions for the same problem and the others as negative examples. To balance the data across problems, we ensure a ratio of 0.15 of positive samples across different languages.
@@
Code Repair: Given a buggy code $C$ in language $L$, generate the corresponding repaired code. We derived the dataset from CodeNet, modifying samples by sequentially inserting, removing, or replacing tokens of different types for fixed ratios for different token types.
Code Generation: Given a prompt $P$ consisting of a function signature and a docstring in language $L$, the task is to generate the function body completion. In the few shot prompting setup, few-shot prompts consisting of three correct functions from the respective MBXP dataset are provided. Few shot examples precede the function completion prompt for each evaluation.
Data sampling.
@@
Our objective is to facilitate an extensive study on transfer across the maximum number of language pairs, which can be helpful for those working with low-resource languages. Sources are selected based on the relative performance of fine-tuned checkpoints compared to a random baseline, in order to determine if there is enough signal in the data to facilitate learning. Languages that demonstrate higher performance than the baseline are included as sources. Given variation in the size of datasets for different languages, we follow the sampling procedure of de Vries et al. (2022) to select the number of training samples and identify potential source or high-resource languages. We first finetune the model with datasets of different number of samples $N$={10K, 30K, 50K, 70K, 100K}. Depending on the number of training examples $N_{L}$ for a language $L$, we randomly upsample languages with We filter the set of source languages for a given task based on the relative performance compared to a baseline model. We select sample sizes of 50K for solution domain classification, 70K for error detection, and 100K for clone detection and code repair.
Models.
@@
Our finetuning experiments are based on CodeT5-base (220M parameters) using the Hugging Face transformers library (Wolf et al., 2020). Code-T5 is an open-source model, pre-trained on eight programming languages (Ruby, JavaScript, Go, Python, Java, PHP, C, and C#). We consider these languages as seen languages and Figure 5 highlights the names of these pre-training languages in red font. Due to its encoder-decoder nature, the model performs well on both code generation and code understanding tasks (Wang et al., 2021) like code repair, error detection, and clone detection, among others. This, along with its relatively small size, makes it a good fit for our empirical study, which requires 58 finetuned models with numerous inference runs each. We keep the same hyperparameters for all the experiments: learning rate of 2e-5, batch size of 8, and the number of epochs set to 20. The few-shot prompting experiment use Llama 3.3 70B-Instruct, [^1] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct an instruction-tuned multilingual LLM (70B parameters) (Grattafiori et al., 2024).
Data Splits
Language Train Test
php 914,924 1,975
rb 912,159 3,963
hs 843,011 2,045
py 785,760 4,235
kt 731,575 1,552
js 713,854 1,600
c 555,924 2,363
scala 473,594 1,705
go 352,181 867
java 266,248 1,338
d 247,954 1,355
cs 233,059 1,151
pl 185,685 1,762
nim 175,839 765
Language Train Test
ml 175,056 2,008
f 172,046 1,253
cpp 150,492 1,792
pas 136,961 1,370
rs 128,271 478
awk 95,312 1,035
jl 88,700 1,737
sh 64,899 1,070
ts 47,414 799
l 45,986 585
swift 38,531 737
fs 31,001 962
scm 22,787 826
cr 21,733 1,356
Language Train Test
sed 21,584 1,349
lua 20,688 1,253
pyx 18,184 931
dc 17,322 1,706
vb 7,003 1,132
octave 5,370 1,476
vim 4,176 745
cob 3,437 1,038
clj 1,058 910
dart 506 480
moon 242 453
ex 330 549
Table 1: Number of train and test samples for Clone Detection for each language.
Train and test splits for Error Detection and Solution Domain Classification are provided from the original dataset as described in the dataset statistics from xCodeEval (Khan et al., 2023). For Code Repair, 50,000 training examples and 1,000 test examples are synthetically generated for each language. For Clone Detection, a distinct set of problems is used for train and test splits, including languages with a minimum of 450 test examples. Table 1 shows the number of examples for Clone Detection for each language. In the few-shot prompting experiment, we conduct a cross-lingual evaluation across all language pairs, using nearly 1,000 examples for each. We additionally use the provided synthetic canonical solutions to measure dataset features across overlapping problem solutions.
4 Results and Discussion
This section addresses the research questions RQ1–RQ3 stated in the introduction.
4.1 Transfer analysis